For the cash flows shown, determine the future worth in year 5.

Year Cash Flow($/year) Estimated i Per Year
0 5000 8%
1–4 6000 8%
5 9000 16%

The future worth in 5 years is $
Answer #1

Answer:

giventhat
FV=PV*(1+i)^n *(1+j)^m

i is for n year and j is for m year rate
and
FV=A*(P|A,i,n)

FV=5000*(1.08^4)*(1.16)+6000*(P|F,i,n)*(1.16)+9000
=5000*(1.08^4)*(1.16)+6000*4.779328*(1.16)+9000
=50154.9579
The future value is $50154.9579